OpenStudy (anonymous):

Carla's plant was 10 2/3 inches tall. Yolanda's was 8 3/4 inches tall. How much taller was Carla's plant than Yolanda's? A. 1 11/12 in B. 2 1/12 in C. 2 5/12 in D. 2 11/12

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

subtract the two

OpenStudy (danjs):

these i would turn to improper fractions first, then put over common denominator

sammixboo (sammixboo):

\(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\) So, this is a problem dealing with subtracting fractions. We have to subtract \(\color{blueviolet}{10\dfrac{2}{3}}~ -~~\color{limegreen}{8\dfrac{3}{4}}\) Do you know the first step in subtracting fractions?

OpenStudy (anonymous):

Finding the Common detonators?

OpenStudy (danjs):

no dont blow them up

sammixboo (sammixboo):

Correct =) Do you know the common multiples of 3 and 4?

OpenStudy (anonymous):

12

sammixboo (sammixboo):

Oops I am sorry. I made a mistake give me a second

sammixboo (sammixboo):

Good, so after we have found the common denominator we should have this \(\color{blueviolet}{10\dfrac{8}{12}}~ -~~\color{limegreen}{8\dfrac{9}{12}}\) Because 12/3 = 4, so \(\color{blueviolet}{10\dfrac{2\times4}{3\times4}}\) And 12/4 = 3, so \(\color{limegreen}{8\dfrac{3\times3}{4\times3}}\) Do you know how to subtract this? \(\color{blueviolet}{10\dfrac{8}{12}}~ -~~\color{limegreen}{8\dfrac{9}{12}}\)

OpenStudy (anonymous):

Not really

sammixboo (sammixboo):

OK so let's make these fractions improper. Do you know how to do that?

OpenStudy (anonymous):

Yep hold on

sammixboo (sammixboo):

OK =)

OpenStudy (anonymous):

128/12 and 105/12

OpenStudy (anonymous):

Hello?

sammixboo (sammixboo):

Correct =) Sorry, so now we have  All you need to do is subtract 128-105 and keep the denominator 12